OpenStudy (anonymous):

Anybody knows how to simplified rational expression : (x^2+2x)/(12x+54) - (3-x)/(8x+36)

OpenStudy (mertsj):

\[\frac{x^2+2x}{12x-54}-\frac{3-x}{8x-36}\]

OpenStudy (mertsj):

1. Change the minus sign to a plus sign and change the numerator of the second fraction to x-3 2. Factor the numerator of the first fraction and the denominators of both fractions. Write both fractions with a common denominator. 3. Add the fractions.

OpenStudy (whpalmer4):

4. simplify again, if possible

OpenStudy (anonymous):

this is what I got: (x^2+7x-9)/ ((6)(2x+9) ..is this correct?

OpenStudy (whpalmer4):

doesn't look like it.

OpenStudy (anonymous):

After finding common denominators I have 2x(x+2)+3(x-3) / 12(2x+9)

OpenStudy (whpalmer4):

hmm, I'm unsure how you are going from 12x-54 and 8x-36 as denominators to 12(2x+9)...

OpenStudy (anonymous):

for the both demonstrators I have 6(2x+9) and 4(2x-9)

OpenStudy (whpalmer4):

\[\frac{x^2+2x}{12x-54} -\frac{3-x}{8x-36} = \frac{x(x+2)}{6(2x-9)} + \frac{x-3}{4(2x-9)}\]

OpenStudy (anonymous):

Yes that what I have!

OpenStudy (anonymous):

and the common denominator is 2(2x-9)

OpenStudy (anonymous):

(2x+9)

OpenStudy (whpalmer4):

you need to be more careful with your signs...no, you just said you have "\(6(2x\color{red}{+}9) \)and \(4(2x-9)\)" no, how are you changing the sign in the denominator?!?

OpenStudy (whpalmer4):

12x-54 8x-36 no plus sign with the constant term in either one!

OpenStudy (anonymous):

I meant 6(2x+9) and 4(2x+9) then common denominato= 2(2x+9)

OpenStudy (whpalmer4):

where are you getting this +?!? the denominators have - signs in them!

OpenStudy (whpalmer4):

12x-54 = 6(2x-9) 8x-36 = 4(2x-9)

OpenStudy (anonymous):

I realized all of them are possitive.. not 12x-54.. instead 12+54 and 8x+36

OpenStudy (whpalmer4):

well, were you planning to tell anyone else?

OpenStudy (whpalmer4):

well, your answer will be simpler this way, at least!

OpenStudy (whpalmer4):

I have to go to dinner. I'll check back later to see how you did. good luck.